### v1

# Rebuilding ROME : Resolving Model Collapse during Sequential Model Editing

Abstract
Recent work on model editing using Rank-One Model Editing (ROME), a popular model editing method, has shown that there are certain facts that the algorithm is unable to edit without breaking the model. Such edits have previously been called disabling edits Gupta et al. (2024). These disabling edits cause immediate model collapse and limits the use of ROME for sequential editing. In this paper, we make two main contributions. Firstly, we show that model collapse with ROME only happens when making edits using the CounterFact dataset Meng et al. (2022a) and does not happen when using the zsRE dataset Levy et al. (2017). Secondly, we find that disabling edits are an artifact of the original implementation of ROME. With this paper, we provide a more stable implementation ROME, which we call r-ROME and show that we no longer observe model collapse when making large scale sequential edits with ROME.
Rebuilding ROME : Resolving Model Collapse during
Sequential Model Editing
Akshat Gupta, Gopala Anumanchipalli UC Berkeley akshat.gupta@berkeley.edu
1 Introduction
Large language models (LLMs) are expensive to train and the knowledge contained in these models gets obsolete with time. Model editing or knowledge editing (Yao et al., 2023) has recently come out as a popular method to update knowledge in large language models (LLMs). Knowledge editing is done in two main ways - either by adding information in-context (Mitchell et al., 2022; Zhong et al., 2023; Cohen et al., 2023), or using methods that modify the parameters of underlying model (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a, b; Tan et al., 2023). In this paper, we focus on one popular parameter-modifying model editing methods called ROME (Rank-One Model Editing) Meng et al. (2022a).
While a lot of model editing approaches perform well when making singular edits, editing multiple facts in a model still remains a challenge for parameter-modifying model editing methods. One way to make multiple edits to the same model is through sequential editing Yao et al. (2023) - where we make a series of single edits to a model by modifying the parameters of the model after every edit. Recent works have started studying the effects of sequential editing and show a gradual decay in model’s abilities as more edits are made to the model (Meng et al., 2022b; Yao et al., 2023; Gu et al., 2024; Gupta et al., 2024). Additionally, it was found that one popular model editing technique called ROME Meng et al. (2022a) was prone to a sudden collapse of the model by a single edit (Gupta et al., 2024; Yang et al., 2024; Hu et al., 2024). This effect was first observed in Gupta et al. (2024) during sequential editing. The collapse included complete loss of downstream performance, inability to recall previously editing facts and loss of the ability to even get edited. Such facts were named disabling edits by Gupta et al. (2024) and were later independently observed by Yang et al. (2024); Hu et al. (2024). Text generation examples for models post-collapse from a disabling edit are shown in Figure 1.
Figure 1: A typical generation example after a disabling edit is compared to a normal model edit using ROME. The bold and underlined part in the text is the prompt given to the model, which is followed by the text generated by the model.
Disabling edits are detrimental for knowledge editing at scale. While a gradual model degradation is expected as we make sequential edits to a model Gupta et al. (2024), disabling edits lead to a sudden model collapse irrespective of when the disabling fact is edited, making sequential editing impossible. An example of this can be seen in Figure 3(a), where instead of allowing gradual model degradation when doing sequential editing like in Figure 4(b), the presence of disabling edits lead to a sudden and immediate model collapse.
(a) CounterFact - ROME
(b) zsRE - ROME
Figure 2: This figure shows distribution of edits along |Delta| and Normalized Entropy metric for edits using ROME on CounterFact and zsRE dataset. These edits are for GPT2-XL model but we see a similar distribution for GPT-J.
In this paper, we aim to find the source of these disabling edits. We first introduce two metrics for identifying disabling edits - generation entropy and the norm of matrix update. We plot edits made by ROME along these two dimensions. We then perform large scale editing of GPT2-XL Radford et al. (2019) and GPT-J Wang and Komatsuzaki (2021) using ROME on two popular model editing datasets - CounterFact Meng et al. (2022a) and zsRE (Levy et al., 2017; Mitchell et al., 2021; Meng et al., 2022a). We find that disabling edits only exist when editing facts from the CounterFact dataset and not the zsRE dataset. After long inquiry into the reasons behind model collapse, we decide to rebuild the ROME code-base ourselves. To our surprise, we find that the problem of model collapse goes away with our implementation of ROME and we are able to perform large-scale sequential model editing using the algorithm. With this paper, we share our new ROME code-base and invite researchers to use it for model editing. Our implementation of ROME, which we call r-ROME, can be found here.
2 Background
The focus of our paper is a model editing algorithm called ROME (Rank-One Model Editing) Meng et al. (2022a). ROME is a parameter-modifying model editing method, which means that it inserts knowledge into LLMs by changing their parameters. ROME falls under the "locate-and-edit" category for model editing methods, where knowledge is updated by first identifying the layers in the model corresponding to stored knowledge and then modifying the weights for the specific layers.
Facts are usually added in a key-value format, where a key is the vector representation of a query-phrase and the value is the vector representation of the target object. For example, when adding a new fact - "The president of USA is John Cena", the query-phrase here is "The president of USA is" and the target object is "John Cena". Editing in ROME is done using a pair of vectors - $(k_{e},v_{e})$ that represent a new fact being added. $k_{e}$, also called the key-vector is a vector representation of the query-phrase, and $v_{e}$, or the value-vector is the vector representation of the target object. The weights of the specific layer being edited in ROME are updated from $W_{0}$ to $\hat{W}$ by inserting a new fact $(k_{e},v_{e})$ using the following equation:
$\hat{W}$ $=W_{0}+\Delta\hskip 10.0pt$ (1)
$\text{where}\hskip 10.0pt\Delta$ $=(v_{e}-W_{0}k_{e})\frac{k_{e}^{T}C_{0}^{-1}}{k_{e}^{T}C_{0}^{-1}k_{e}}$
where $\Delta$ is the update to the current weight matrix being edited such that the new fact represented by $(k_{e},v_{e})$ is incorporated.
Dataset Implementation Efficacy Generalization Locality Fluency Score
ES $\uparrow$ EM $\uparrow$ PS $\uparrow$ PM $\uparrow$ NS $\uparrow$ NM $\uparrow$ GE $\uparrow$ S $\uparrow$
ROME Original $99.92$ $99.68$ $96.29$ $71.58$ $75.8$ $10.25$ $621.96$ $89.32$
Ours $99.78$ $99.64$ $97.4$ $71.16$ $74.86$ $10.0$ $621.08$ $89.16$
Table 1: A comparison between the original implementation Meng et al. (2022a) and our implementation of ROME on standard model editing metrics for GPT2-XL. The above table is created using a random sample of 5000 edits from the CounterFact dataset. We find that our implementation performs equally as well as the original implementation of ROME.
2.1 Evaluating Model Editing
Model editing is usually evaluated along three metrics - reliability, generalization and locality. Reliability measures if a fact was successfully added in a model, generalization measures if the edited fact is recalled through paraphrases of the prompt used to edit the fact, whereas locality measures if editing of one fact affects other facts stored inside a model. We follow standard model editing metrics proposed in Meng et al. (2022a). We refer the reader to Yao et al. (2023); Meng et al. (2022a) for a more comprehensive review of model editing metrics. Additionally, we also evaluated the model on downstream task performance as proposed by (Gu et al., 2024; Gupta et al., 2024), which becomes especially important when making sequential edits to the same model. We follow Gupta et al. (2024) and evaluate the edited model on four tasks from the GLUE Wang et al. (2018) benchmark - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019) for doing downstream evaluation.
3 Experiments
3.1 Metrics to Identify Disabling Edits
Disabling edits Gupta et al. (2024) are defined as singular knowledge edits made to the model that renders it useless, leading to sudden loss of ability to do downstream tasks or any kind of meaningful generation. Gupta et al. (2024) also showed one way of identifying disabling edits was the unusually large norm of the update matrix. In other words, $|\Delta|$ in equation 1 was unusually higher when compared to the $|\Delta|$ for normal edits. We will use a modified version of $|\Delta|$ as the first metric to identify disabling edits, where we normalize $|\Delta|$ by the number of elements in the matrix.
Additionally, Figure 1 shows a typical example of model collapse where the model constantly repeats a single word. This repeated word is most often the edited answer inserted into the model. The simplest metric to identify such a model collapse is to calculate the entropy over the probability distribution of vocabulary elements of text generated from the model. For this, a probability distribution is calculated over the vocabulary of a sample generation consisting of ten generations, and is normalized by the vocabulary size to remove the effect of the size of vocabulary. This constrains the entropy between 0 and 1. If the model collapses as shown in Figure 1, we expected the normalized entropy to be small and concentrated around a handful of words.
Note that this metric is a simplification of the generation entropy metric proposed by Meng et al. (2022a), where they take a weighted average of the entropy over of vocabulary of bigrams and trigrams. We chose to calculate entropy over just the unigram vocabulary words as it is cheaper and does the job, while normalization helps comparing entropy values for generations of different vocabulary size.
(a) Original Implementation
(b) Our Implementation
Figure 3: This figure shows the difference between the original updates and our implementation for GPT2-XL on the CounterFact dataset for 5k edits. As can be seen, our implementation significantly reduces potential disabling edits, with $|\Delta|$ being way lower.
3.2 Searching for Disabling Edits
The first set of experiments we do is to search for disabling edits. We do this by making large-scale singular model edits using ROME on GPT2-XL, on two popular model editing datasets - CounterFact Meng et al. (2022a) and zsRE (Levy et al., 2017; Mitchell et al., 2021). We measure the above mentioned metrics as shown in Figure 2.
We make the following two observations. Firstly, we observe that edits made using the zsRE dataset are very consistent along the two metrics, which means that the norm of the update are consistently small and generation entropy consistently high. The text generations made by all post-edit models when editing zsRE are consistently good and coherent, and do not have the repetitiveness found in disabling edits. We thus conclude that disabling edits are not present when editing facts using the zsRE dataset. Secondly, we see two clusters when editing facts from the CounterFact dataset. We find that certain edits have larger values of $|\Delta|$ for ROME, indicating the presence of disabling edits as seen in (Gupta et al., 2024; Yang et al., 2024; Hu et al., 2024). This shows that edits made using the CounterFact dataset can lead to model collapse. The above observation is quite surprising. It shows that there is a fundamental difference in the updates made to the model when editing facts using zsRE dataset and the CounterFact dataset. Model updates made using zsRE dataset have smaller $|\Delta|$ values and lead to healthy model behavior post editing.
(a) Downstream Evaluation
(b) $|\Delta|$
Figure 4: Sequential editing using original implementation of ROME on GPT-J (6B).
(a) Downstream Evaluation
(b) $|\Delta|$
Figure 5: Sequential editing with our re-implementation of ROME on GPT-J (6B).
The underlying reason for the difference in behavior of post-edit model is unclear, but is likely related to the characteristics of the two datasets. zsRE and CounterFact dataset differ in three major ways. Firstly, the CounterFact dataset contains counterfactual facts. This means that this dataset edits the model in such a way that lower probability objects are being inserted into the model, whereas in zsRE, we are inserting factually correct facts. The second difference in the two datasets is that zsRE edits facts using a question-answering prompt, whereas CounterFact prompts the model in a text completion format. This means that the key-vector in zsRE for the fact mentioned in section 2 corresponds to a question like "Who is the president of USA?" whereas for CounterFact it corresponds to "The president of USA is". Thirdly, all facts in CounterFact are one-word facts, are largely also tokenized into a single token for GPT2-XL and GPT-J, whereas most facts in zsRE contain multiple words. The difference in edits made by the two datasets is possibly due to one of the underlying reasons, but it is hard to answer this question without eliminating the confounding factors. Such a study is beyond the scope of this paper, and hence we continue to focus on removing disabling edits during model editing with ROME.
3.3 Fixing Disabling Edits
After a long inquiry into the optimization objective of ROME and studying its original code-base and different properties of the edits made by the algorithm, we found no reason for the $|\Delta|$ of certain edits to be so large. As a last resort, we re-implemented ROME ourselves. To our absolute surprise, we found that our re-implementation of ROME does not lead to disabling edits. The first evidence of this can be seen in Figure 3, where the $|\Delta|$ of the updates are orders of magnitude smaller for our implementation when compared to the original implementation. The comparison of original Meng et al. (2022a) and our implementation of ROME on standard model editing metrics is shown in Table 1. This table shows that our implementation performs equally well as original implementation of ROME, while having much smaller $|\Delta|$ values. We make no change to the update equation of ROME and re-implement it as is.
3.4 Sequential Editing with ROME
A large part of the reason why disabling edits went unnoticed initially was because large scale edits were not performed on the same model and the model was not evaluated on downstream tasks. Recent works (Gupta et al., 2024; Yang et al., 2024; Hu et al., 2024) observe disabling edits as a result performing sequential edits. Thus a final litmus test of our re-implementation of ROME is to perform sequential editing. The results for sequential editing can be seen in Figures 4 and 5.
Figure 4 shows a typical case of sequential editing using the original ROME code-base for GPT-J, where the presence of a disabling edit leads to large $|\Delta|$ and leads to model collapse, as can be seen by an immediate loss of downstream performance in Figure 3(a). With our re-implementation of ROME (Figure 5), we see that $|\Delta|$ is orders of magnitude smaller which allows the model to maintain its general abilities and avoids model collapse. This enables large scale sequential model editing without loss of performance. Similar trends can be seen for GPT2-XL in Figures 6 and 7.
4 Conclusion
In this paper, we show that model edits made using the original implementation of ROME lead to unstable model edits eventually causing model collapse. Our re-implementation of ROME, which we call r-rome (code) prevents model collapse and leads to stable and scalable model edits, thus making sequential editing possible using ROME.
Appendix A Sequential Editing on GPT2-XL
(a) Downstream Evaluation
(b) $|\Delta|$
Figure 6: Sequential editing using original implementation of ROME on GPT2-XL (1.5B).
(a) Downstream Evaluation
(b) $|\Delta|$
Figure 7: Sequential editing with our re-implementation of ROME on GPT2-XL (1.5B).